In a hearing test, subjects estimate the loudness (in decibels) of sound, and the results are; 12, 68, 74, 66, 37, 52, 71, 90, 65, 76, 73, 22. What are the outer fences?
Answer #1

The values can br re written as

12, 22, 37, 52, 65, 66, 68, 71, 73, 74, 76, 90

Hence five point summary of the data will be as

Min= 12

Q1 =44.5

Q2= 67

Q3= 73.5

Max= 90

Hence IQR = Q3- Q1

=73.5-44.5

= 29

Hence the Outer fence is calculated as

Upper fence = Q3 + 1.5* IQR

= 73.5+ 1.5* 29

=73.5+ 43.5

=117

Lower Fence

= Q1- 1.5 *IQR

=44.5- 43.5

=1